OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (cp9454):

B

OpenStudy (anonymous):

It would have to be B

OpenStudy (anonymous):

The negatives of the points cancel out in the equation y1-y2=m(x1-x2).

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

B becuz I plugged it into the equation y-y1=m(x-x1) y-(-1)=4(x-(3)= y+1=4x+3 it would be B show work guys!

OpenStudy (anonymous):

i explained myself!